Confer a value to be significantly low if it’s z score less than or equal to -2 or consider a value to be significantly high if it’s z score is greater than or equal to 2.

a data set lists weights(grams) of a type of a coin. Those weights have a mean of 5.74712g and a standard deviation of 0.06038g. Identify the weights that are significantly low or significantly high


1
Expert's answer
2021-09-29T00:47:59-0400

SOLUTION

The z-score is given by: "Z-score=(x-\mu)/\sigma"

Where

"x=score\\\mu=mean=5.74712g\\\sigma=stdev=0.06038"

Given that the z-score is -2, then the lowest cost will be:

"-2=\frac{(x-5.74712)}{0.06038}"

Solving for x we get:

"-0.12076=x-5.74712\\5.74712-0.12076=x\\x=5.62636"

ANSWER: The lowest score is 5.62636

Given that the z-score is 2, then the lowest cost will be:

"2=\frac{(x-5.74712)}{0.06038}"

Solving for x we get:

"0.12076=x-5.74712\\5.74712+0.12076=x\\x=5.86788"

ANSWER: The highest score is 5.86788